OpenStudy (anonymous):

Whats 100cm = how many inches rounded to the nearest tenth

OpenStudy (unklerhaukus):

\[2.45[\text{cm}]=1[\text{in}]\\1=\frac{1[\text{in}]}{2.45[\text{cm}]}\] \[100[\text{cm}]\\=100[\text{cm}]\times1\\=100[\text{cm}]\times\frac{1[\text{in}]}{2.45[\text{cm}]}\\=\frac{100}{2.45}[\text{in}]\]